OpenStudy (anonymous):

In which situation is the acceleration of the car negative? A) The velocity of a car was 75 km/h over 4 hours. B) The velocity of a car reduced from 50 km/h to 35 km/h over one minute. C) The velocity of a car increased from 40 km/h to 75 km/h over 15 minutes. D) The velocity of a car was 45 km/h at 2:00 pm and at 4:00 pm the velocity of the car was 85 km/h.

OpenStudy (anonymous):

@Brrandyn Help please!! Im begging you!

OpenStudy (brrandyn):

Acceleration is defined as the rate of velocity over a time interval: \[a =\frac{ Δv }{ t}\] Therefore, the difference in the final velocity must be less than the initial velocity. \[a = \frac{ Vf - Vi }{ t }\] The answer is: B) 35 - 50 = negative acceleration.

OpenStudy (anonymous):

Thank you so much!!!!!! ;D

OpenStudy (brrandyn):

Sure! :) Sorry for the slow typing.

OpenStudy (anonymous):

Its fine! :)